OpenStudy (anonymous):

Convert 0.3110M of BSA to mg/mL BSA: 66.5kDa/mole = 1.104 x 10^-19 gm/mole

OpenStudy (aaronq):

This is incorrect "66.5kDa/mole = 1.104 x 10^-19 gm/mole" 1 Da = 1 g/mol, so 66.5 kDa= 66500 g/mol \(M=\dfrac{n_{solute}}{L_{solution}}=\dfrac{m_{solute}}{M^{solute}_r*L_{solution}}\) 1 L = 1000 mL \(m_{solute}=M*M^{solute}_r*L_{solution}=0.3110~M *66500~g/mol*1~L=20681.5~g\) 20681500 mg in 1000 mL concentration = 20681.5 mg/mL which seems stupidly high, but hey